# Coupled Distributional Random Expert Distillation for World Model Online Imitation Learning

### v1

@@ 4.2 DMControl Experiments
We conduct experiments on 6 tasks in DMControl [43] environments. For low-dimensional tasks, we utilize 100 expert trajectories, while for high-dimensional tasks, we use 500 expert trajectories. Details on environment dimensionality can be found in Appendix C. Our CDRED model performs comparably to IQ-MPC on the Hopper Hop, Walker Run, and Humanoid Walk tasks. However, in Cheetah Run and Dog Stand, IQ-MPC experiences long-term instability, causing the agent to fail after extensive online training. On the Reacher Hard task, IQ-MPC struggles with an overly powerful discriminator, which prevents it from learning an expert-level policy. The model-free methods in baseline algorithms fail to achieve stable, expert-level performance on these tasks. The episode reward results are shown in Figure 5.
@@ 4.3 Vision-based Experiments
In addition to experiments using state-based observations, we also benchmark our method on tasks with visual observations. Specifically, we select three tasks from DMControl [43] with visual observations. To create these visual datasets, we render visual observations based on state-based expert trajectories, replacing the original state-based observations in the expert data. For each task, we use 100 expert trajectories generated by a trained TD-MPC2 model [26]. We show our results in Figure 6. Interestingly, we observe that visual IQ-MPC encounters an issue with an overly powerful discriminator in the Cheetah Run task when using trajectories generated by a trained state-based TD-MPC2 policy, where state observations are replaced by RGB images rendered from those states. However, IQ-MPC performs well when using expert trajectories generated by a TD-MPC2 policy trained directly on visual observations.
@@ A.1 Architectural Details
We show the overall model architecture via a Pytorch style notation. We leverage layernorm [1] and Mish activations [35] for our model. The detialed architecture is displayed as following:
@@ A.2 Hyperparameter Details
The remaining hyperparameters are consistent with those used in TD-MPC2 [26].
@@ B.1 Training Algorithm
0:  $\theta,\theta^{-}$: randomly initialized network parameters          $\eta,\tau,\lambda,\mathcal{B}_{\pi},\mathcal{B}_{E}$: learning rate, soft update coefficient, horizon discount coefficient, behavioral buffer, expert buffer
@@
[…]
@@
[…]
@@
        for $t=0...H$ do
[…]
@@
[…]
@@ B.2 Planning Algorithm
0:  $\theta:$ learned network parameters          $\mu^{0},\sigma^{0}$: initial parameters for $\mathcal{N}$          $N,N_{\pi}$: number of sample/policy trajectories          $\mathbf{s}_{t},H$: current state, rollout horizon
@@
3:     Sample $N$ trajectories of length $H$ from $\mathcal{N}(\mu^{j-1},(\sigma^{j-1})^{2}\mathrm{I})$
@@
6:        for step $t=0..H-1$ do
7:           $\mathbf{z}_{t+1}\leftarrow d_{\theta}(\mathbf{z}_{t},\mathbf{a}_{t})$vartriangleleft Latent transition
@@
9:           $\phi_{\Gamma}=\phi_{\Gamma}+\gamma^{t}R_{\theta}(\mathbf{z}_{t},\mathbf{a}_{t})$vartriangleleft Estimate reward with CDRED reward model
@@
11:        $\phi_{\Gamma}=\phi_{\Gamma}+\gamma^{H}Q_{\theta}(\mathbf{z}_{H},\mathbf{a}_{H})$vartriangleleft Terminal Q value
@@ Appendix C Task Details and Environment Specifications
We consider 12 continuous control tasks in locomotion control and robot manipulation. We leverage 6 manipulation tasks in Meta-World [52], 6 locomotion tasks in DMControl [43] and 3 tasks in ManiSKill2 [18]. In this section, we list the environment specifications for completeness in Table 2, Table 3 and Table 4.
@@ D.1 Experiments on ManiSkill2
We further evaluate our method on additional manipulation tasks in ManiSkill2 [18], achieving stable and competitive results on the pick cube, lift cube, and turn faucet tasks. Notably, IQL+SAC [16] and IQ-MPC [34] also perform relatively well in these scenarios. Table 5 summarizes the success rates of each method across the ManiSkill2 tasks.
@@
Pick Cube 0.61$\pm$0.13 0.00$\pm$0.00 0.79$\pm$0.05 0.87$\pm$0.04
Lift Cube 0.85 $\pm$ 0.04 0.01$\pm$0.01 0.89$\pm$0.02 0.93$\pm$0.03
Turn Faucet 0.82$\pm$0.04 0.00$\pm$0.00 0.73$\pm$0.08 0.84$\pm$0.08
@@ D.3 Additional Comparison with HyPE
Hybrid IRL [39] is a recently proposed method for performing inverse reinforcement learning and imitation learning using hybrid data. In this section, we compare our approach with the model-free method (HyPE) introduced in their work. Our method achieves superior empirical performance on three DMControl locomotion tasks, including the high-dimensional Humanoid Walk task. The results are presented in Figure 10.
Figure 10: Comparison with HyPE We compare our CDRED approach with the HyPE method [39] on the Hopper Hop, Cheetah Run, and Humanoid Walk tasks. Among these, the Humanoid Walk task is high-dimensional, while the others are low-dimensional. Our approach demonstrates superior empirical performance and improved sampling efficiency on these tasks.
@@ D.4 Additional Comparison with SAIL
Support-weighted Adversarial Imitation Learning (SAIL) [46] is an extension of Generative Adversarial Imitation Learning (GAIL) [27] that enhances performance by integrating Random Expert Distillation (RED) rewards [45]. In this section, we present an additional comparative analysis between our proposed CDRED method and SAIL. The experimental results are illustrated in Figure 11.
Figure 11: Comparison with SAIL We compare our CDRED approach with the SAIL method [46] on the Walker Run, Cheetah Run, and Humanoid Walk tasks. Among these, the Humanoid Walk task is high-dimensional, while the others are low-dimensional. SAIL fails to learn in the high-dimensional Humanoid Walk task while our approach achieves nearly expert-level performance. Overall, our approach demonstrates superior empirical performance and improved sampling efficiency on these tasks.
@@ D.5 Robustness Analysis under Noisy Dynamics
We conduct an additional analysis to evaluate the robustness of our model under noisy environment dynamics. Following the evaluation protocol of Hybrid IRL [39], we introduce noise by adding a trembling probability, $p_{\text{tremble}}$. During interactions with the environment, the agent executes a random action with probability $p_{\text{tremble}}$ and follows the action generated by the policy for the remaining time. Our empirical results demonstrate that our model exhibits robustness to noisy dynamics, as its performance only slightly deteriorates from the expert level when noise is introduced. The results for the Cheetah Run and Walker Run tasks are presented in Figure 12.
@@ D.6 Quantitative Analysis of Training Stability
To assess the training stability of our algorithm, we examine the mean and maximum gradient norms throughout the training process. This approach is similar to the analysis conducted in TD-MPC2 [26]. We compare the gradient norms of our method with those of IQ-MPC [34], a world model online imitation learning approach that employs an adversarial formulation, on DMControl tasks. Our results indicate that the gradient norms of our approach are significantly smaller than those of IQ-MPC, suggesting superior training stability. The detailed comparison is presented in Table 6.
@@ D.7 Advantages Compared to Current Methods Involving Adversarial Training
The current existing methods [34, 32, 37, 51] for world model online imitation learning often involve adversarial training, following the similar problem formulation as GAIL [27] or IQ-Learn [16]. IQ-MPC [34] adopted inverse soft-Q objective for critic learning while CMIL [32], V-MAIL [37] and EfficientImitate [51] leveraged GAIL style reward modeling. In terms of IQ-Learn, an improved version of GAIL, although its policy can be computed by applying a softmax to the Q-value in discrete control, effectively converting a min-max problem into a single maximization [16], it still requires the maximum entropy RL objective for policy updates in continuous control settings. In such cases, IQ-Learn performs adversarial training between the policy and the critic, which leads to stability issues similar to those encountered in GAIL. IQ-MPC, while performing well in various complex scenarios such as high-dimensional locomotion control and dexterous hand manipulation, still encounters challenges in some cases. These challenges include an imbalance between the discriminator and the policy, as well as long-term instability. These issues stem from using an adversarially trained Q-function as the critic. While IQ-MPC attempts to mitigate them by incorporating regularization terms during the training process, it doesn’t fully resolve the problem. Figure 13 illustrates the drawbacks of IQ-MPC in some cases, namely an overly powerful discriminator and long-term instability. We also demonstrate the quantitative results for training stability analysis in Appendix D.6.
@@
Figure 13: Drawbacks of Methods Including Adversarial Training We demonstrate the drawbacks of IQ-MPC [34] in some tasks, which employs adversarial training for online imitation learning. An overly powerful discriminator (Left) leads to sub-optimal policy learning, while long-term instability (Right) of adversarial training prevents IQ-MPC from maintaining expert-level performance during extended online training. Our CDRED method, which replaces adversarial training with density estimation, is immune to these issues.
@@
The generative adversarial training process is often prone to instability [19]. IQ-MPC employs generative adversarial training between the policy and the critic, and it also encounters this challenge. To mitigate this issue, IQ-MPC leverages gradient penalty from [19] to enforce Lipschitz condition of the gradients in a form of:
@@
In the gradient penalty, $(\mathbf{\hat{s}}_{t},\mathbf{\hat{a}}_{t})$ are data points on straight lines between expert and behavioral distributions, which are generated by linear interpolation. Although it counters the problem to some extent, the performance of IQ-MPC is still not satisfactory in some tasks such as Reacher in DMControl and Meta-World robotics manipulation tasks, for which we will refer to our experimental results in Section 4. An overly powerful discriminator often causes the Q-value difference between the policy and expert distributions to diverge, as noted by [34]. Specifically, this divergence is reflected in the gap between the expected Q-values under the expert distribution, $\mathbb{E}_{(\mathbf{s},\mathbf{a})_{(0:H)}\sim\mathcal{B}_{E}}Q(\mathbf{z}_{t},\mathbf{a}_{t})$, and the policy distribution, $\mathbb{E}_{(\mathbf{s},\mathbf{a})_{(0:H)}\sim\mathcal{B}_{\pi}}Q(\mathbf{z}_{t},\mathbf{a}_{t})$. While IQ-MPC can mitigate this divergence to some extent through gradient penalty, it does not eliminate the difference entirely, indicating that the policy does not achieve expert-level performance. We show the Q difference plot in a problematic case in Figure 14.
@@ D.8 Improvement of Constructing the Reward Model on the Latent Space
Original RND [4] and Random Expert Distillation [45] train their reward or bonus models directly on the original observation space. In contrast, we found that constructing the CDRED reward model using the latent representations from a world model yields better empirical performance. This highlights the superior properties of latent representations, which enable more accurate reward estimation. Furthermore, by training a latent dynamics model within this space, the representations become more dynamics-aware, facilitating the construction of a reward model that effectively captures the underlying dynamics.
@@ Appendix E Proof of Lemma 1
For completeness, we adapt the proof from [49] to construct the proof of Lemma 1. For a latent state-action pair $(\mathbf{z},\mathbf{a})$ sampled from a latent state-action distribution $\rho$. We denote the moments of the distribution of random variable $c(\mathbf{z},\mathbf{a})$ as:
@@ Appendix F Related Works
Recent advancements in imitation learning (IL) have leveraged deep neural networks and diverse methodologies to enhance performance. Generative Adversarial Imitation Learning (GAIL) [27] laid the foundation for adversarial reward learning by formulating it as a min-max optimization problem inspired by Generative Adversarial Networks (GANs) [17]. Several approaches have built on GAIL. Model-based Adversarial Imitation Learning (MAIL) [2] extended GAIL with a forward model trained via data-driven methods. ValueDICE [33] transformed the adversarial framework by focusing on off-policy learning through distribution ratio estimation.
Offline imitation learning has seen significant advancements through approaches like Diffusion Policy [6], which applied diffusion models for behavioral cloning, and Ditto [9], which combined Dreamer V2 [23] with adversarial techniques. Implicit BC [14] demonstrated that supervised policy learning with implicit models improves empirical performance in robotic tasks. DMIL [54] leveraged a discriminator to assess dynamics accuracy and the suboptimality of model rollouts against expert demonstrations in offline IL.
Other innovations focused on integrating advanced reinforcement learning techniques. Inverse Soft Q-Learning (IQ-Learn) [16] reformulated GAIL’s learning objectives, applying them to soft actor-critic [20] and soft Q-learning agents. SQIL [38] contributed an online imitation learning algorithm utilizing soft Q-functions. CFIL [15] introduced a coupled flow method for simultaneous reward generation and policy learning from expert demonstrations. Random Expert Distillation (RED) [45] proposed an alternative method for constructing reward models by estimating the support of the expert policy distribution.
Model-based methods have also played a pivotal role in advancing IL. V-MAIL [37] employed variational models to facilitate imitation learning, while CMIL [32] utilized conservative world models for image-based manipulation tasks. Prior works [11, 28, 30] highlighted the potential of model-based imitation learning in real-world robotics control and autonomous driving. A model-based inverse reinforcement learning approach by Das et al. [8] explored key-point prediction to improve performance in imitation tasks. Hybrid Inverse Reinforcement Learning [39] offered a novel strategy blending online and expert demonstrations, enhancing agent robustness in stochastic settings. EfficientImitate [51] fused EfficientZero [50] with adversarial imitation learning, achieving impressive performance on DMControl tasks [43].
@@
Recent advancements in model-based reinforcement learning (MBRL) utilize learned dynamics models, constructed via data-driven methodologies, to enhance agent learning and decision-making. MBPO [31] introduced a model-based policy optimization algorithm that ensures stepwise monotonic improvement. Extending this to offline RL, MOPO [53] incorporated a penalty term in the reward function based on the uncertainty of the dynamics model to manage distributional shifts effectively. MBVE [13] augmented model-free agents with model-based rollouts to improve value estimation.
Many approaches focus on constructing dynamics models in latent spaces. PlaNet [22] pioneered this direction by proposing a recurrent state-space model (RSSM) with an evidence lower bound (ELBO) training objective, addressing challenges in partially observed Markov decision processes (POMDPs). Building on PlaNet, the Dreamer algorithms [21, 23, 24] leveraged learned world models to simulate future trajectories in a latent space, enabling efficient learning and planning. The TD-MPC series [25, 26] further refined latent-space modeling by developing a scalable world model for model predictive control, utilizing a temporal-difference learning objective to improve performance. Similarly, MuZero [40] combined a latent dynamics model with tree-based search to achieve strong performance in discrete control tasks, blending planning and policy learning seamlessly. The EfficientZero series [50, 47] enhances MuZero, achieving superior sampling efficiency in visual reinforcement learning tasks.

### v2

@@ 4.2 DMControl Experiments
We conduct experiments on 6 tasks in DMControl [tassa2018deepmind] environments. For low-dimensional tasks, we utilize 100 expert trajectories, while for high-dimensional tasks, we use 500 expert trajectories. Details on environment dimensionality can be found in Appendix C. Our CDRED model performs comparably to IQ-MPC on the Hopper Hop, Walker Run, and Humanoid Walk tasks. However, in Cheetah Run and Dog Stand, IQ-MPC experiences long-term instability, causing the agent to fail after extensive online training. On the Reacher Hard task, IQ-MPC struggles with an overly powerful discriminator, which prevents it from learning an expert-level policy. The model-free methods in baseline algorithms fail to achieve stable, expert-level performance on these tasks. The episode reward results are shown in Figure 5.
@@ 4.3 Vision-based Experiments
In addition to experiments using state-based observations, we also benchmark our method on tasks with visual observations. Specifically, we select three tasks from DMControl [tassa2018deepmind] with visual observations. To create these visual datasets, we render visual observations based on state-based expert trajectories, replacing the original state-based observations in the expert data. For each task, we use 100 expert trajectories generated by a trained TD-MPC2 model [hansen2023td]. We show our results in Figure 6. Interestingly, we observe that visual IQ-MPC encounters an issue with an overly powerful discriminator in the Cheetah Run task when using trajectories generated by a trained state-based TD-MPC2 policy, where state observations are replaced by RGB images rendered from those states. However, IQ-MPC performs well when using expert trajectories generated by a TD-MPC2 policy trained directly on visual observations.
@@ A.1 Architectural Details
We show the overall model architecture via a Pytorch style notation. We leverage layernorm [ba2016layer] and Mish activations [misra2019mish] for our model. The detialed architecture is displayed as following:
@@ A.2 Hyperparameter Details
The remaining hyperparameters are consistent with those used in TD-MPC2 [hansen2023td].
@@ B.1 Training Algorithm
[…]
@@
   $\mathcal{B}_{\pi}\leftarrow\mathcal{B}_{\pi}\cup(\mathbf{s}_{t},\mathbf{a}_{t},r_{t},\mathbf{s}^{\prime}_{t})\vartriangleleft$ Add to behavioral buffer
@@
   $(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}^{\prime}_{t})_{0:H}\sim\mathcal{B}_{\pi}\cup\mathcal{B}_{E}\vartriangleleft$ Combine behavioral and expert batch
   $\mathbf{z}_{0}=h_{\theta}(\mathbf{s}_{0})\vartriangleleft$ Encode first observation
@@
   for t=0Ht=0...H do
    $\mathbf{z}_{t+1}=d_{\theta}(\mathbf{z}_{t},\mathbf{a}_{t})\vartriangleleft$ Unrolling using the latent dynamics model
    $\hat{q}_{t}=Q(\mathbf{z}_{t},\mathbf{a}_{t})\vartriangleleft$ Estimate the Q value
    $\mathbf{z}^{\prime}_{t}=h(\mathbf{s}^{\prime}_{t})\vartriangleleft$ Encode the ground-truth next state
    $\hat{r}_{t}=R(\mathbf{z}_{t},\mathbf{a}_{t})\vartriangleleft$ Estimate Reward using the CDRED reward model
    $q_{t}=\hat{r}_{t}+\gamma Q(\mathbf{z}^{\prime}_{t},\pi(\mathbf{z}^{\prime}_{t}))\vartriangleleft$ Compute the TD target using the estimated reward
@@
   Compute model loss $\mathcal{L}\vartriangleleft$ Equation 15
   Compute policy prior loss $\mathcal{L}^{\pi}\vartriangleleft$ Equation 16
   $\theta\leftarrow\theta-\frac{1}{H}\eta\nabla_{\theta}(\mathcal{L}+\mathcal{L}^{\pi})\vartriangleleft$ Update online network
   $\theta^{-}\leftarrow(1-\tau)\theta^{-}+\tau\theta\vartriangleleft$ Soft update
@@ B.2 Planning Algorithm
[…]
@@
3:  Sample NN trajectories of length $H$ from $\mathcal{N}(\mu^{j-1},(\sigma^{j-1})^{2}\mathrm{I})$
@@
6:   for step t=0..H1t=0..H-1 do
[…]
@@
[…]
@@
[…]
@@ Appendix C Task Details and Environment Specifications
We consider 12 continuous control tasks in locomotion control and robot manipulation. We leverage 6 manipulation tasks in Meta-World [yu2020meta], 6 locomotion tasks in DMControl [tassa2018deepmind] and 3 tasks in ManiSKill2 [gu2023maniskill2]. In this section, we list the environment specifications for completeness in Table 2, Table 3 and Table 4.
@@ D.1 Experiments on ManiSkill2
We further evaluate our method on additional manipulation tasks in ManiSkill2 [gu2023maniskill2], achieving stable and competitive results on the pick cube, lift cube, and turn faucet tasks. Notably, IQL+SAC [garg2021iq] and IQ-MPC [li2024rewardfreeworldmodelsonline] also perform relatively well in these scenarios. Table 5 summarizes the success rates of each method across the ManiSkill2 tasks.
@@
[…]
@@ D.3 Additional Comparison with HyPE
Hybrid IRL [ren2024hybrid] is a recently proposed method for performing inverse reinforcement learning and imitation learning using hybrid data. In this section, we compare our approach with the model-free method (HyPE) introduced in their work. Our method achieves superior empirical performance on three DMControl locomotion tasks, including the high-dimensional Humanoid Walk task. The results are presented in Figure 10.
Figure 10: Comparison with HyPE We compare our CDRED approach with the HyPE method [ren2024hybrid] on the Hopper Hop, Cheetah Run, and Humanoid Walk tasks. Among these, the Humanoid Walk task is high-dimensional, while the others are low-dimensional. Our approach demonstrates superior empirical performance and improved sampling efficiency on these tasks.
@@ D.4 Additional Comparison with SAIL
Support-weighted Adversarial Imitation Learning (SAIL) [SAIL] is an extension of Generative Adversarial Imitation Learning (GAIL) [ho2016generative] that enhances performance by integrating Random Expert Distillation (RED) rewards [wang2019random]. In this section, we present an additional comparative analysis between our proposed CDRED method and SAIL. The experimental results are illustrated in Figure 11.
Figure 11: Comparison with SAIL We compare our CDRED approach with the SAIL method [SAIL] on the Walker Run, Cheetah Run, and Humanoid Walk tasks. Among these, the Humanoid Walk task is high-dimensional, while the others are low-dimensional. SAIL fails to learn in the high-dimensional Humanoid Walk task while our approach achieves nearly expert-level performance. Overall, our approach demonstrates superior empirical performance and improved sampling efficiency on these tasks.
@@ D.5 Robustness Analysis under Noisy Dynamics
We conduct an additional analysis to evaluate the robustness of our model under noisy environment dynamics. Following the evaluation protocol of Hybrid IRL [ren2024hybrid], we introduce noise by adding a trembling probability, $p_{\text{tremble}}$. During interactions with the environment, the agent executes a random action with probability $p_{\text{tremble}}$ and follows the action generated by the policy for the remaining time. Our empirical results demonstrate that our model exhibits robustness to noisy dynamics, as its performance only slightly deteriorates from the expert level when noise is introduced. The results for the Cheetah Run and Walker Run tasks are presented in Figure 12.
@@ D.6 Quantitative Analysis of Training Stability
To assess the training stability of our algorithm, we examine the mean and maximum gradient norms throughout the training process. This approach is similar to the analysis conducted in TD-MPC2 [hansen2023td]. We compare the gradient norms of our method with those of IQ-MPC [li2024rewardfreeworldmodelsonline], a world model online imitation learning approach that employs an adversarial formulation, on DMControl tasks. Our results indicate that the gradient norms of our approach are significantly smaller than those of IQ-MPC, suggesting superior training stability. The detailed comparison is presented in Table 6.
@@ D.7 Advantages Compared to Current Methods Involving Adversarial Training
The current existing methods [li2024rewardfreeworldmodelsonline, kolev2024efficient, rafailov2021visual, yin2022planning] for world model online imitation learning often involve adversarial training, following the similar problem formulation as GAIL [ho2016generative] or IQ-Learn [garg2021iq]. IQ-MPC [li2024rewardfreeworldmodelsonline] adopted inverse soft-Q objective for critic learning while CMIL [kolev2024efficient], V-MAIL [rafailov2021visual] and EfficientImitate [yin2022planning] leveraged GAIL style reward modeling. In terms of IQ-Learn, an improved version of GAIL, although its policy can be computed by applying a softmax to the Q-value in discrete control, effectively converting a min-max problem into a single maximization [garg2021iq], it still requires the maximum entropy RL objective for policy updates in continuous control settings. In such cases, IQ-Learn performs adversarial training between the policy and the critic, which leads to stability issues similar to those encountered in GAIL. IQ-MPC, while performing well in various complex scenarios such as high-dimensional locomotion control and dexterous hand manipulation, still encounters challenges in some cases. These challenges include an imbalance between the discriminator and the policy, as well as long-term instability. These issues stem from using an adversarially trained Q-function as the critic. While IQ-MPC attempts to mitigate them by incorporating regularization terms during the training process, it doesn’t fully resolve the problem. Figure 13 illustrates the drawbacks of IQ-MPC in some cases, namely an overly powerful discriminator and long-term instability. We also demonstrate the quantitative results for training stability analysis in Appendix D.6.
@@
Figure 13: Drawbacks of Methods Including Adversarial Training We demonstrate the drawbacks of IQ-MPC [li2024rewardfreeworldmodelsonline] in some tasks, which employs adversarial training for online imitation learning. An overly powerful discriminator (Left) leads to sub-optimal policy learning, while long-term instability (Right) of adversarial training prevents IQ-MPC from maintaining expert-level performance during extended online training. Our CDRED method, which replaces adversarial training with density estimation, is immune to these issues.
@@
The generative adversarial training process is often prone to instability [gulrajani2017improved]. IQ-MPC employs generative adversarial training between the policy and the critic, and it also encounters this challenge. To mitigate this issue, IQ-MPC leverages gradient penalty from [gulrajani2017improved] to enforce Lipschitz condition of the gradients in a form of:
@@
In the gradient penalty, $(\mathbf{\hat{s}}_{t},\mathbf{\hat{a}}_{t})$ are data points on straight lines between expert and behavioral distributions, which are generated by linear interpolation. Although it counters the problem to some extent, the performance of IQ-MPC is still not satisfactory in some tasks such as Reacher in DMControl and Meta-World robotics manipulation tasks, for which we will refer to our experimental results in Section 4. An overly powerful discriminator often causes the Q-value difference between the policy and expert distributions to diverge, as noted by [li2024rewardfreeworldmodelsonline]. Specifically, this divergence is reflected in the gap between the expected Q-values under the expert distribution, $\mathbb{E}_{(\mathbf{s},\mathbf{a})_{(0:H)}\sim\mathcal{B}_{E}}Q(\mathbf{z}_{t},\mathbf{a}_{t})$, and the policy distribution, $\mathbb{E}_{(\mathbf{s},\mathbf{a})_{(0:H)}\sim\mathcal{B}_{\pi}}Q(\mathbf{z}_{t},\mathbf{a}_{t})$. While IQ-MPC can mitigate this divergence to some extent through gradient penalty, it does not eliminate the difference entirely, indicating that the policy does not achieve expert-level performance. We show the Q difference plot in a problematic case in Figure 14.
@@ D.8 Improvement of Constructing the Reward Model on the Latent Space
Original RND [burda2018exploration] and Random Expert Distillation [wang2019random] train their reward or bonus models directly on the original observation space. In contrast, we found that constructing the CDRED reward model using the latent representations from a world model yields better empirical performance. This highlights the superior properties of latent representations, which enable more accurate reward estimation. Furthermore, by training a latent dynamics model within this space, the representations become more dynamics-aware, facilitating the construction of a reward model that effectively captures the underlying dynamics.
@@ Appendix E Proof of Lemma 1
For completeness, we adapt the proof from [yang2024exploration] to construct the proof of Lemma 1. For a latent state-action pair $(\mathbf{z},\mathbf{a})$ sampled from a latent state-action distribution $\rho$. We denote the moments of the distribution of random variable $c(\mathbf{z},\mathbf{a})$ as:
@@ Appendix F Related Works
Recent advancements in imitation learning (IL) have leveraged deep neural networks and diverse methodologies to enhance performance. Generative Adversarial Imitation Learning (GAIL) [ho2016generative] laid the foundation for adversarial reward learning by formulating it as a min-max optimization problem inspired by Generative Adversarial Networks (GANs) [goodfellow2014generative]. Several approaches have built on GAIL. Model-based Adversarial Imitation Learning (MAIL) [baram2016model] extended GAIL with a forward model trained via data-driven methods. ValueDICE [kostrikov2019imitation] transformed the adversarial framework by focusing on off-policy learning through distribution ratio estimation.
Offline imitation learning has seen significant advancements through approaches like Diffusion Policy [chi2023diffusionpolicy], which applied diffusion models for behavioral cloning, and Ditto [demoss2023ditto], which combined Dreamer V2 [hafner2020mastering] with adversarial techniques. Implicit BC [florence2022implicit] demonstrated that supervised policy learning with implicit models improves empirical performance in robotic tasks. DMIL [zhang2023discriminator] leveraged a discriminator to assess dynamics accuracy and the suboptimality of model rollouts against expert demonstrations in offline IL.
Other innovations focused on integrating advanced reinforcement learning techniques. Inverse Soft Q-Learning (IQ-Learn) [garg2021iq] reformulated GAIL’s learning objectives, applying them to soft actor-critic [haarnoja2018soft] and soft Q-learning agents. SQIL [reddy2019sqil] contributed an online imitation learning algorithm utilizing soft Q-functions. CFIL [freund2023coupled] introduced a coupled flow method for simultaneous reward generation and policy learning from expert demonstrations. Random Expert Distillation (RED) [wang2019random] proposed an alternative method for constructing reward models by estimating the support of the expert policy distribution.
Model-based methods have also played a pivotal role in advancing IL. V-MAIL [rafailov2021visual] employed variational models to facilitate imitation learning, while CMIL [kolev2024efficient] utilized conservative world models for image-based manipulation tasks. Prior works [englert13model, hu2022model, igl2022symphony] highlighted the potential of model-based imitation learning in real-world robotics control and autonomous driving. A model-based inverse reinforcement learning approach by das2021model explored key-point prediction to improve performance in imitation tasks. Hybrid Inverse Reinforcement Learning [ren2024hybrid] offered a novel strategy blending online and expert demonstrations, enhancing agent robustness in stochastic settings. EfficientImitate [yin2022planning] fused EfficientZero [ye2021mastering] with adversarial imitation learning, achieving impressive performance on DMControl tasks [tassa2018deepmind].
@@
Recent advancements in model-based reinforcement learning (MBRL) utilize learned dynamics models, constructed via data-driven methodologies, to enhance agent learning and decision-making. MBPO [janner2019trust] introduced a model-based policy optimization algorithm that ensures stepwise monotonic improvement. Extending this to offline RL, MOPO [yu2020mopo] incorporated a penalty term in the reward function based on the uncertainty of the dynamics model to manage distributional shifts effectively. MBVE [feinberg2018model] augmented model-free agents with model-based rollouts to improve value estimation.
Many approaches focus on constructing dynamics models in latent spaces. PlaNet [hafner2019learning] pioneered this direction by proposing a recurrent state-space model (RSSM) with an evidence lower bound (ELBO) training objective, addressing challenges in partially observed Markov decision processes (POMDPs). Building on PlaNet, the Dreamer algorithms [hafner2019dream, hafner2020mastering, hafner2023mastering] leveraged learned world models to simulate future trajectories in a latent space, enabling efficient learning and planning. The TD-MPC series [hansen2022temporal, hansen2023td] further refined latent-space modeling by developing a scalable world model for model predictive control, utilizing a temporal-difference learning objective to improve performance. Similarly, MuZero [schrittwieser2020mastering] combined a latent dynamics model with tree-based search to achieve strong performance in discrete control tasks, blending planning and policy learning seamlessly. The EfficientZero series [ye2021mastering, wang2024efficientzero] enhances MuZero, achieving superior sampling efficiency in visual reinforcement learning tasks.